# Quantum Steenrod operations of symplectic resolutions

### v2

@@ 3.4. Beyond Springer resolutions
Examples which satisfy 3.23 include the Springer resolution $X=T^{*}(G/B)$ and its parabolic generalizations $T^{*}(G/P)$, and quiver vareities of type $A$.
The stable envelopes, satisfying the properties 3.7, 3.8, exist in a much wider generality of conical symplectic resolutions, see [MO19, Section 3] and [SZ20]. In particular, for a choice of a generic cocharacter $\beta:S^{1}\to T$ (and a choice of a polarization data, see [MO19, Section 3.3.2]) one obtains the corresponding stable envelope map
@@
Under 3.23, 3.8 applies and the stable envelope classes $\mathrm{Stab}_{\beta}(x)\in H^{\dim_{\mathbb{C}}X}_{\mathbf{T}}(X)$ for isolated fixed points $x\in X^{T}$ form a basis in $H^{*}_{\mathbf{T}}(X)_{\mathrm{loc}}$. Moreover, the dual basis is given by the opposite stable envelopes $\mathrm{Stab}_{-\beta}$.
@@
Corollary 3.25 (cf. 3.15, 3.19).
Let $X$ be a conical symplectic resolution satisfying 3.23. Then the equivariant quantum Steenrod operations $\Sigma_{b}^{\mathbf{T}}$ and $\mathbb{S}_{\beta}$ for nonnegative $\beta:S^{1}\to{\mathbf{T}}$ may be defined on $H^{*}_{\mathbf{T}}(X)_{\mathrm{loc}}$.
@@
As in the case of 3.11, 3.24 is sufficient to show that the curves in moduli space of holomorphic curves with incidence constraints on stable envelope cycles and their duals are contained in a compact subset of $X$. Using the proper, holomorphic projection to the affine base $X\to X_{0}$, one can proceed as in the proof of 3.14 and 3.18 to show the regularity of the moduli spaces of perturbed solutions. The nonnegativity condition on cocharacter from 3.16 modifies so that the cocharacter pairs nonnegatively with ${\mathbf{T}}$-weights of the ${\mathbf{T}}$-representation $H^{0}(X;\mathcal{O})$. ∎
@@ 4. $p$-curvature of equivariant quantum connections
In this section, we review the equivariant quantum connections of conical symplectic resolutions from [BMO11], [MO19]. The equivariant quantum connections of symplectic resolutions are special in that (i) they admit lifts to the integers $\mathbb{Z}$ (or with bounded denominators) in a suitable basis, (ii) dependence on Novikov parameters is a rational function (a priori, it is only a power series) and (ii) they admit a compatible collection of intertwining operators given by the shift operators. Such connections fall into the general class of flat connections studied in [EV24b, EV24a], in which structural results about their $p$-curvature is proven.
@@ 4.1. Equivariant quantum connections
Recall that $\lambda,h$ represent the equivariant parameters for the action of torus $T\leq{\mathbf{T}}$ on $X$ (preserving the holomorphic symplectic form), and the action of $S^{1}\leq{\mathbf{T}}$ on $X$ (scaling the holomorphic symplectic form), respectively. We consider $H^{*}_{\mathbf{T}}(X)_{\mathrm{loc}}$ as a $H^{*}_{\mathbf{T}}(\mathrm{pt})_{\mathrm{loc}}$-module with basis given by the stable envelope classes $\mathrm{Stab}_{+}(w)$ indexed by Weyl group elements $w\in W$, see 3.7. This basis of (localized) equivariant cohomology of $X$ is referred to as the stable basis.
@@
Consider the equivariant quantum connection of the Springer resolution $X=T^{*}(G/B)$. The connection satisfies the following properties:
@@
    The quantum connection commutes with the shift operators $\mathbb{S}_{\beta}$ for every choice of $\beta$.
@@
Let $b\in H^{2}(X;\mathbb{Z})$ identified with some weight of the maximal torus $T$. Denote by $R$, $R^{\vee}$ the set of roots and coroots and identify positive coroots (after a choice of a dominant chamber) $\alpha^{\vee}\in R_{+}^{\vee}\subseteq H_{2}(X;\mathbb{Z})$ with curve degree classes. The formula from [BMO11, Theorem 1.1] characterizes the quantum multiplication operator as
(4.1) $b\ \ast_{\ {\mathbf{T}}}=b\smile_{\mathbf{T}}+\ h\sum_{\alpha^{\vee}\in R^{\vee}_{+}}(b,\alpha^{\vee})\frac{q^{\alpha^{\vee}}}{1-q^{\alpha^{\vee}}}([s_{\alpha}]-1),$
@@
Property (ii) is evident from (4.1). Property (i) follows from the from the computation of the action of $[s_{\alpha}]$ in stable basis [Su17, Lemma 3.2] which we reproduce below:
@@
(4.4) $b\ \ast_{\ {\mathbf{T}}}=b\smile_{{\mathbf{T}}}+\ b\ \ast_{\mathrm{quantum},{\mathbf{T}}}.$
Since all basis elements of the stable basis lie in the same cohomological degree $\dim_{\mathbb{C}}X$ (see 3.8), all entries of the matrix of $b\ \ast_{\ {\mathbf{T}}}$ must be expressions of cohomological degree $2$, depending polynomially on the equivariant parameters $h,\lambda$ of the ${\mathbf{T}}$-action of $X$. Expanding in $h$, one can write in the stable basis $b\smile_{{\mathbf{T}}}=hB^{(0)}_{\mathrm{cl}}+B^{(1)}_{\mathrm{cl}}$ and $b\ \ast_{\mathrm{quantum},{\mathbf{T}}}=hB_{q}^{(0)}$. Note that all entries of the purely quantum part is divisible by $h$, which is evident from (4.1) and is a consequence of the geometric fact that non-$S^{1}$-equivariant Gromov–Witten invariants of $X$ vanish. ∎
@@ 4.2. $p$-curvature
Fix $b\in H^{2}(X;\mathbb{F}_{p})$ a mod $p$ reduction of an integral class, and consider the corresponding equivariant quantum connection $\nabla_{b}=t\partial_{b}+b\ \ast_{{\mathbf{T}}}:H^{*}_{\mathbf{T}}(X;\Lambda){[\![t,\theta]\!]}\to H^{*}_{\mathbf{T}}(X;\Lambda){[\![t,\theta]\!]}$, see 2.19. The $p$-curvature in the direction of $b$ is the degree $2p$ endomorphism of $H^{*}_{\mathbf{T}}(X;\Lambda){[\![t,\theta]\!]}$ defined as
@@
    $F_{b}(b_{0})|_{q^{A}=0}=\mathrm{St}^{\mathbf{T}}(b)\smile_{{\mathbf{T}}}b_{0}$, where $\mathrm{St}^{\mathbf{T}}(b)\in H^{*}_{{\mathbf{T}}}(X;\mathbb{F}_{p}){[\![t,\theta]\!]}$ is the class obtained by the power operation construction as in (2.24) in ${\mathbf{T}}$-equivariant cohomology.
@@
Property (i) follows from that $\nabla_{b}|_{(t,\theta)=0}=b\ \ast_{\mathbf{T}}$. Property (ii) follows from (i) the computation of the classical Steenrod operation $\mathrm{St}(b)=b^{\smile_{\mathbf{T}}p}-t^{p-1}b$ for degree $2$ classes applied to $b\in H^{2}_{\mathbf{T}}(X;\mathbb{F}_{p})$, and (ii) the observation that $\nabla_{b}|_{q^{A}=0}=t\partial_{b}+b\smile_{\mathbf{T}}$ and that $\partial_{b}$ commutes with the classical multiplication $b\smile_{\mathbf{T}}$, hence $\nabla_{b}^{p}=t^{p}\partial_{b}+b^{\smile_{\mathbf{T}}p}$. Property (iii) follows from the fact that $\nabla_{b}$ commutes with the shift operators, 2.28. ∎
@@
Let $\nabla$ be a flat connection (possibly with poles) defined over a sublocus $H_{\mathrm{reg}}:=H\setminus\bigcup_{\alpha}\{q^{\alpha}=1\}$ on a trivial bundle $V\times H_{\mathrm{reg}}\to H_{\mathrm{reg}}$, which depends on a parameter $\hbar\in k$, i.e.
@@
The ($S^{1}$)-equivariant quantum connection of the Springer resolution after the identification of parameters $h/t=\hbar$ forms a periodic pencil of flat connections, considered as a trigonometric connection, by the results of [BMO11]. The bare shift operators $S_{\beta}$ provide the required gauge equivalences; see [BMO11, Proposition 6.1] and 2.28.
@@
[EV24a] Let the spectrum of the connection matrix of the connection $\nabla_{i}$ in a periodic pencil be given by $\Lambda_{ij}(q)$, $j=1,\dots,\mathrm{rank}(V)$. Then the spectrum of the $p$-curvature $F_{i}$ are explicitly given by
@@
In particular, if the connection matrix of $\nabla_{i}$ has simple spectrum, then so does the $p$-curvature $F_{i}$.
In [Lee23, Section 7], we have observed that for the quantum Steenrod operation of $T^{*}\mathbb{P}^{1}$, the specialization of equivariant parameters $\Sigma_{b}|_{h/t=\hbar}$ for $\hbar\in\mathbb{F}_{p}$ annihilates Varchenko’s mod $p$ flat sections of the quantum connection. 4.8 is consonant with this observation; specializing $\hbar\in\mathbb{F}_{p}$ annihilates all eigenvalues.

### v3

@@ 3.4. Beyond Springer resolutions
Examples which satisfy 3.26 include the Springer resolution $X=T^{*}(G/B)$ and its parabolic generalizations $T^{*}(G/P)$, and quiver vareities of type $A$.
The stable envelopes, satisfying the properties 3.8, 3.9, exist in a much wider generality of conical symplectic resolutions, see [MO19, Section 3] and [SZ20]. In particular, for a choice of a generic cocharacter $\beta:S^{1}\to T$ (and a choice of a polarization data, see [MO19, Section 3.3.2]) one obtains the corresponding stable envelope map
@@
Under 3.26, 3.9 applies and the stable envelope classes $\mathrm{Stab}_{\beta}(x)\in H^{\dim_{\mathbb{C}}X}_{\mathbf{T}}(X)$ for isolated fixed points $x\in X^{T}$ form a basis in $H^{*}_{\mathbf{T}}(X)_{\mathrm{loc}}$. Moreover, the dual basis is given by the opposite stable envelopes $\mathrm{Stab}_{-\beta}$.
@@
Corollary 3.28 (cf. 3.18, 3.22).
Let $X$ be a conical symplectic resolution satisfying 3.26. Then the equivariant quantum Steenrod operations $\Sigma_{b}^{\mathbf{T}}$ and $\mathbb{S}_{\beta}$ for nonnegative $\beta:S^{1}\to\mathbf{T}$ may be defined on $H^{*}_{\mathbf{T}}(X)_{\mathrm{loc}}$.
@@
As in the case of 3.13, 3.27 is sufficient to show that the curves in moduli space of holomorphic curves with incidence constraints on stable envelope cycles and their duals are contained in a compact subset of $X$. Using the proper, holomorphic projection to the affine base $X\to X_{0}$, one can proceed as in the proof of 3.16 and 3.21 to show the regularity of the moduli spaces of perturbed solutions. The nonnegativity condition on cocharacter from 3.19 modifies so that the cocharacter pairs nonnegatively with $\mathbf{T}$-weights of the $\mathbf{T}$-representation $H^{0}(X;\mathcal{O})$. ∎
@@ 4. $p$-curvature of equivariant quantum connections
In this section, we review the equivariant quantum connections of conical symplectic resolutions from [BMO11], [MO19]. The equivariant quantum connections of symplectic resolutions are special in that (i) they admit lifts to the integers $\mathbb{Z}$ (or with bounded denominators) in a suitable basis, (ii) dependence on Novikov parameters is a rational function (a priori, it is only a power series) and (iii) they admit a compatible collection of intertwining operators given by the shift operators. Such connections fall into the general class of flat connections studied in [EV24b, EV24a], in which structural results about their $p$-curvature is proven.
@@ 4.1. Equivariant quantum connections
Recall that $\lambda,h$ represent the equivariant parameters for the action of torus $T\leq\mathbf{T}$ on $X$ (preserving the holomorphic symplectic form), and the action of $S^{1}\leq\mathbf{T}$ on $X$ (scaling the holomorphic symplectic form), respectively. We consider $H^{*}_{\mathbf{T}}(X)_{\mathrm{loc}}$ as a $H^{*}_{\mathbf{T}}(\mathrm{pt})_{\mathrm{loc}}$-module with basis given by the stable envelope classes $\mathrm{Stab}_{+}(w)$ indexed by Weyl group elements $w\in W$, see 3.8. This basis of (localized) equivariant cohomology of $X$ is referred to as the stable basis.
@@
Consider the equivariant quantum connection operators $\nabla_{b}:H^{*}_{\mathbb{Z}/p\times\mathbf{T}}(X;\Lambda)\to H^{*}_{\mathbb{Z}/p\times\mathbf{T}}(X;\Lambda)$, indexed by $b\in H^{2}_{\mathbf{T}}(X;\mathbb{Z})$, of the Springer resolution $X=T^{*}(G/B)$. The connection satisfies the following properties:
@@
    The quantum connection commutes with the shift operators $\mathbb{S}_{\beta}:H^{*}_{\mathbb{Z}/p\times\mathbf{T}}(X;\Lambda)\to H^{*}_{\mathbb{Z}/p\times\mathbf{T}}(X;\Lambda)$ for every choice of $\beta$.
@@
By linearity of quantum product with respect to equivariant parameters, it suffices to consider $b\in H^{2}(X;\mathbb{Z})$. Let $b\in H^{2}(X;\mathbb{Z})$ be identified with some weight of the maximal torus $T$. Denote by $R$, $R^{\vee}$ the set of roots and coroots and identify positive coroots (after a choice of a dominant chamber) $\alpha^{\vee}\in R_{+}^{\vee}\subseteq H_{2}(X;\mathbb{Z})$ with curve degree classes. The formula from [BMO11, Theorem 1.1] characterizes the quantum multiplication operator as
(4.1) $b\ \ast_{\mathbf{T}}=b\cup_{\mathbf{T}}+\ h\sum_{\alpha^{\vee}\in R^{\vee}_{+}}(b,\alpha^{\vee})\frac{q^{\alpha^{\vee}}}{1-q^{\alpha^{\vee}}}([s_{\alpha}]-1),$
@@
Property (ii) is evident from (4.1). Property (i) follows from the computation of the action of $[s_{\alpha}]$ in stable basis [Su17, Lemma 3.2] which we reproduce below:
@@
Example 4.2.
For $X=T^{*}\mathbb{P}^{1}$, the non-equivariant cohomology $H^{*}(X)$ is of rank $2$; the equivariant cohomology $H^{*}_{\mathbf{T}}(X)$ has two stable basis elements of cohomological degree $2$ given by $\mathrm{Stab}_{+}(1)$ and $\mathrm{Stab}_{+}(s)$ indexed by the Weyl group $W=\{1,s\}$. The action of the Weyl group is given as in Eq. 4.2.
The precise form of the connection (in $S^{1}$-equivariant cohomology, where we forget the $T=S^{1}$-action) in the stable basis is given in [Lee23, Lemma 7.1].
@@
(4.4) $b\ \ast_{\mathbf{T}}=b\cup_{\mathbf{T}}+\ b\ast_{\mathrm{quantum},\mathbf{T}}.$
Since all basis elements of the stable basis lie in the same cohomological degree $\dim_{\mathbb{C}}X$ (see 3.9), and the quantum product $b\ast_{\mathbf{T}}$ is an operator of degree $2$, all entries of the matrix of $b\ \ast_{\mathbf{T}}$ must be expressions of cohomological degree $2$. These expressions depend polynomially on the equivariant parameters $h,\lambda$ of the $\mathbf{T}$-action of $X$; the regularity of moduli spaces ensure that the equivariant parameters appear with non-negative powers, and the Novikov parameter $q^{A}$ has degree zero (as $c_{1}=0$). Expanding in $h$, one can write in the stable basis $b\cup_{\mathbf{T}}=hB^{(0)}_{\mathrm{cl}}+B^{(1)}_{\mathrm{cl}}$ and $b\ \ast_{\mathrm{quantum},\mathbf{T}}=hB_{q}^{(0)}$. Note that all entries of the purely quantum part is divisible by $h$, which is evident from (4.1) and is a consequence of the geometric fact that non-$S^{1}$-equivariant Gromov–Witten invariants of $X$ vanish. ∎
@@ 4.2. $p$-curvature
Fix $b\in H^{2}_{\mathbf{T}}(X;\mathbb{F}_{p})$ a mod $p$ reduction of an integral class, and consider the corresponding equivariant quantum connection $\nabla_{b}=t\partial_{b}+b\ \ast_{\mathbf{T}}:H^{*}_{\mathbf{T}}(X;\Lambda){[\![t,\theta]\!]}\to H^{*}_{\mathbf{T}}(X;\Lambda){[\![t,\theta]\!]}$, see 2.19. The $p$-curvature in the direction of $b$ is the degree $2p$ endomorphism of $H^{*}_{\mathbf{T}}(X;\Lambda){[\![t,\theta]\!]}$ defined as
@@
    $F_{b}(b_{0})|_{q^{A}=0}=\mathrm{St}^{\mathbf{T}}(b)\cup_{\mathbf{T}}b_{0}$, where $\mathrm{St}^{\mathbf{T}}(b)\in H^{*}_{\mathbf{T}}(X;\mathbb{F}_{p}){[\![t,\theta]\!]}$ is the class obtained by the power operation construction as in (2.24) in $\mathbf{T}$-equivariant cohomology.
@@
Property (i) follows from that $\nabla_{b}|_{(t,\theta)=0}=b\ \ast_{\mathbf{T}}$. Property (ii) follows from the definition of the $p$-curvature $F_{b}=\nabla_{b}^{p}-t^{p-1}\nabla_{b}$ together with the following two observations. (a) The computation of the classical Steenrod operations (denoted $P^{0}$ and $P^{1}$) gives
(4.7) $\mathrm{St}^{\mathbf{T}}(b)=P^{1}(b)-t^{p-1}P^{0}(b)=b^{\cup_{\mathbf{T}}p}-t^{p-1}b$
for degree $2$ classes applied to $b\in H^{2}_{\mathbf{T}}(X;\mathbb{F}_{p})$ (the Bockstein vanishes as we assume that $H^{*}(X)$ is torsion-free). The notation $b^{\cup_{\mathbf{T}}p}$ is the $p$-fold cup power of $b$ in $\mathbf{T}$-equivariant cohomology. (b) The connection satisfies $\nabla_{b}|_{q^{A}=0}=t\partial_{b}+b\cup_{\mathbf{T}}$ and that $\partial_{b}$ commutes with the classical multiplication $b\cup_{\mathbf{T}}$, hence $\nabla_{b}^{p}|_{q^{A}=0}=t^{p}\partial_{b}+b^{\cup_{\mathbf{T}}p}$. Property (iii) follows from the fact that $\nabla_{b}$ commutes with the shift operators, 2.28. ∎
@@
Let $\nabla$ be a flat connection (possibly with poles) defined over a sublocus $H_{\mathrm{reg}}:=H\setminus\bigcup_{\alpha}\{q^{\alpha}=1\}$ on a trivial bundle $V\times H_{\mathrm{reg}}\to H_{\mathrm{reg}}$, which depends on a parameter $\hbar\in k$ linearly, i.e.
@@
The ($S^{1}$)-equivariant quantum connection of the Springer resolution after the identification of parameters $h/t=\hbar$ forms a periodic pencil of flat connections, considered as a trigonometric connection, by the results of [BMO11]. Namely, the quantum product has only a single power of $h$; we divide by $t$ as the quantum connection is only a connection in the usual convention after dividing by $t$. The bare shift operators $S_{\beta}$ provide the required gauge equivalences; see [BMO11, Proposition 6.1] and 2.28.
@@
[EV24a, Theorem 3.3] Let the spectrum of the connection matrix of the connection $\nabla_{i}$ in a periodic pencil be given by $\Lambda_{ij}(q)$, $j=1,\dots,\mathrm{rank}(V)$. Then the spectrum of the $p$-curvature $F_{i}$ are explicitly given by
@@
In particular, if the connection matrix of $\nabla_{i}$ has simple spectrum (distinct eigenvalues of multiplicity one), then so does the $p$-curvature $F_{i}$.
In [Lee23, Section 7], we have observed that for the quantum Steenrod operation of $T^{*}\mathbb{P}^{1}$, the specialization of equivariant parameters $\Sigma_{b}|_{h/t=\hbar}$ for $\hbar\in\mathbb{F}_{p}$ annihilates Varchenko’s mod $p$ flat sections of the quantum connection. 4.9 is consonant with this observation; specializing $\hbar\in\mathbb{F}_{p}$ annihilates all eigenvalues.
@@ 5. $p$-curvature and quantum Steenrod operations
5.1. A table
For the proof of the main theorem below, we use several properties of the (equivariant) quantum Steenrod operations, the $p$-curvature of the (equivariant) quantum connection, and the shift operators. Here we record, for the convenience of the reader, their linearity properties and respective limits under different specializations of equivariant parameters. We denote by $q^{A}$ the Novikov parameters, $(t,\theta)$ the $\mathbb{Z}/p$-equivariant parameters, $(\lambda,h)$ the $\mathbf{T}=T\times S^{1}$-parameters.
operator $q\mapsto 0$ $(t,\theta)\mapsto 0$ $h\mapsto 0$ commutativity with $\nabla_{a}$ linear over
$\Sigma_{b}^{\mathbf{T}}$ $\mathrm{St}^{\mathbf{T}}(b)\cup_{\mathbf{T}}$ $b^{\ast_{\mathbf{T}}p}\ast_{\mathbf{T}}$ $\mathrm{St}^{T}(b)\cup_{T}$ $[\Sigma_{b}^{\mathbf{T}},\nabla_{a}]=0$ $q,t,\theta,h,\lambda$
$F_{b}$ $\mathrm{St}^{\mathbf{T}}(b)\cup_{\mathbf{T}}$ $b^{\ast_{\mathbf{T}}p}\ast_{\mathbf{T}}$ $\mathrm{St}^{T}(b)\cup_{T}$ $[F_{b},\nabla_{a}]=0$ $q,t,\theta,h,\lambda$
$\mathbb{S}_{\beta}=\Phi_{\beta}\circ S_{\beta}$ twist Seidel map $[\mathbb{S}_{\beta},\nabla_{a}]=0$ $q,t,\theta$
For the first and the second row, the corresponding claims can be found in Section 2.3 and Section 4.2.
The limit $h\mapsto 0$ for $\Sigma_{b}^{\mathbf{T}}$ and $F_{b}$ being $\mathrm{St}^{T}(b)\cup_{T}$ is specific to the target being a holomorphic symplectic manifold. A priori, the limit $h\mapsto 0$ of each operator corresponds to the $T$-equivariant (as opposed to $\mathbf{T}$-equivariant) quantum Steenrod operation $\Sigma_{b}^{T}$, and the $p$-curvature of the $T$-equivariant quantum connection $(t\partial_{b}+b\ast_{T})^{p}-t^{p-1}(t\partial_{b}+b\ast_{T})$ (as opposed to $F_{b}=(t\partial_{b}+b\ast_{\mathbf{T}})^{p}-t^{p-1}(t\partial_{b}+b\ast_{\mathbf{T}})$), respectively. If we work non-equivariantly with respect to the dilating $S^{1}$-action, the Gromov–Witten theory is trivial (see [BMO11, Section 1.3] for our target $X$, hence setting $h=0$ has the effect of eliminating all $q\neq 0$ terms in the respective operators, reducing the limit to the first column $q\mapsto 0$.
For the last row, “twist” refers to the reparametrization morphism which sends $(t,\theta,\lambda,h)\mapsto(t,\theta,\lambda+\beta_{\lambda}t,h+\beta_{h}t)$ for the cocharacter $\beta=(\beta_{\lambda},\beta_{h})\in\mathrm{Lie}(T)\times\mathrm{Lie}(S^{1})=\mathrm{Lie}(\mathbf{T})$. The shift operator is only defined for $X$ with a torus action, hence the corresponding entry is empty. When we work only $\mathbf{T}$-equivariantly (as opposed to $\mathbb{Z}/p\times\mathbf{T}$-equivariantly) in its construction, the shift operators reduce to the Seidel representation [LJ20, Iri17, Sei97].